OpenStudy (anonymous):

7/9 divided by 11/15?? simplidied form

OpenStudy (anonymous):

=7/9 * 15/11 = 7/3 * 5/11 = 35/33 or 1 2/35

OpenStudy (anonymous):

it says 1 2/35 is incorrect

OpenStudy (anonymous):

ok ill take another look

OpenStudy (anonymous):

i've done it on my calculator to check - it comes up with 35/33 which is 1 2/33 - sorry i made an error converting its 1 2/33

OpenStudy (anonymous):

much appericated

OpenStudy (anonymous):

yw